OpenStudy (anonymous):

Compute f(x)=2t(sin(t^2))dt from -1 to x

OpenStudy (anonymous):

t^2 = u use this :)

OpenStudy (anonymous):

ok, thanks :) i was substituting sint^2 by mistake

OpenStudy (anonymous):

is it cosx^2-cos1 or cos1-cosx^2?

OpenStudy (anonymous):

\[2t*\sin t^2 dt\] \[t^2= u\] \[dt = \frac{dy}{2t}\] \[ \frac{2t}{2t}\sin u*du\]

OpenStudy (anonymous):

-cos u - cos t^2 -cos x^2 + cos 2

OpenStudy (anonymous):

sorry cos 1 - cos x^2